cassieerobyn:

Shown below are the steps a student took to solve the simple interest formula A = P(1 + rt) for r.

Vocaloid:

Shown below are the steps a student took to solve the simple interest formula A = P(1 + rt) for r. A = P(1 + rt) A = P+ rt A - P= rt r = (A - P)/t Explain the algebraic error in the student’s work. Then, state how the equation can be modified to produce the correct result.

Vocaloid:

Thoughts? As a major hint there’s an error in the distribution of P across the parentheses.